# Functional Central Limit Theorem for the simultaneous subgraph count of dynamic Erdős-Rényi random graphs

### v1

@@ 4. Main result
$\eta_{i,g,N}(s,t)$ $=\mathbb{I}\{g{\,\sqsubset\,}G_{N}(t)\}-p^{\mathcal{E}({H}_{i})}(t)-\mathbb{I}\{g{\,\sqsubset\,}G_{N}(s)\}+p^{\mathcal{E}({H}_{i})}(s).$ (8.3)
@@
$x_{k}=\begin{cases}s,\qquad k\in\{1,2\},\\ r,\qquad k\in\{3,4\},\end{cases}\qquad y_{k}=\begin{cases}t,\qquad k\in\{1,2\},\\ s,\qquad k\in\{3,4\}.\end{cases}\qquad i_{k}=\begin{cases}i,\qquad k\in\{1,2\},\\ j,\qquad k\in\{3,4\}.\end{cases}$
@@
$\Delta_{N,i,j}(r,s,t)=\frac{\mathcal{F}_{N}^{2}({H}_{i})\mathcal{F}_{N}^{2}({H}_{j})}{\mathcal{F}^{\rm opt}_{N}({H}_{i})\mathcal{F}^{\rm opt}_{N}({H}_{j})}\sum_{\boldsymbol{g}\in{{\tt G}_{N}(i,j)}}\mathbb{E}\left\{\prod_{k=1}^{4}\eta_{i_{k},g_{k},N}(x_{k},y_{k})\right\}$ (8.6)
@@
$\Omega_{k}$ =$\bigl{\{}\mathbb{I}\{g_{k}{\,\sqsubset\,}G_{N}(y_{k})\}\not=\mathbb{I}\{g_{k}{\,\sqsubset\,}G_{N}(x_{k})\}\bigr{\}}.$ (8.7)
@@
Continuation of proof of Lemma 5.5. Combining the bounds presented in Lemma 8.1 with (8.8) and the evident bound $\eta_{i_{k},g_{k},N}(x_{k},y_{k})\leqslant 2$, we obtain that
$\mathbb{E}\left\{\prod_{k=1}^{4}\eta_{i_{k},g_{k},N}(x_{k},y_{k})\right\}$
$\leqslant$ $\>\mathbb{P}\left\{\Omega_{1}^{c}\Omega_{2}^{c}\Omega_{3}^{c}\Omega_{4}^{c}\right\}\prod_{k=1}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k}})}\mathcal{E}({H}_{i_{k}})p_{+}^{\mathcal{E}({H}_{i_{k}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\>+$
$\frac{2}{3!}\sum_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}\mathbb{P}\left\{\Omega_{k_{1}}\Omega_{k_{2}}^{c}\Omega_{k_{3}}^{c}\Omega_{k_{4}}^{c}\right\}\prod_{q=2}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k_{q}}})}\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\>+$
$\frac{4}{2!2!}\sum_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}\mathbb{P}\left\{\Omega_{k_{1}}\Omega_{k_{2}}\Omega_{k_{3}}^{c}\Omega_{k_{4}}^{c}\right\}\prod_{q=3}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k_{q}}})}\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\>+$
@@
$\leqslant$ $\>\prod_{k=1}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k}})}\mathcal{E}({H}_{i_{k}})p_{+}^{\mathcal{E}({H}_{i_{k}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\>+$ (8.9)
$\frac{2}{3!}\sum_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}C_{0}\varrho_{N}^{\mathcal{E}(g_{k_{1}})}\left\lvert t-r\right\rvert\prod_{q=2}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k_{q}}})}\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\>+$
$\frac{4}{2!2!}\sum_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}C_{0}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}})}\left\lvert t-r\right\rvert\prod_{q=3}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k_{q}}})}\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\>+$
@@
$16\,C_{0}\varrho_{N}^{\mathcal{E}(g_{1}\cup g_{2}\cup g_{3}\cup g_{4})}\left\lvert t-r\right\rvert^{2}$
@@
$\prod_{k=1}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k}})}\mathcal{E}({H}_{i_{k}})p_{+}^{\mathcal{E}({H}_{i_{k}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\leqslant C_{1}F_{1}(N)\left\lvert t-r\right\rvert^{2}$ (8.10)
@@
$\frac{2}{3!}\sum_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}C_{0}\varrho_{N}^{\mathcal{E}(g_{k_{1}})}\left\lvert t-r\right\rvert\prod_{q=2}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k_{q}}})}\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\leqslant C_{2}F_{2}(N)\left\lvert t-r\right\rvert^{2}$ (8.11)
@@
$\frac{4}{2!2!}\sum_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}C_{0}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}})}\left\lvert t-r\right\rvert\prod_{q=3}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k_{q}}})}\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\leqslant 4!\,C_{3}F_{3}(N)\left\lvert t-r\right\rvert^{2}$ (8.12)
@@
$\frac{8}{3!}\sum_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}C_{0}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}})}\left\lvert t-r\right\rvert^{2}\varrho_{N}^{\mathcal{E}({H}_{i_{k_{4}}})}\mathcal{E}({H}_{i_{k_{4}}})p_{+}^{\mathcal{E}({H}_{i_{k_{4}}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\leqslant 4!\,C_{4}F_{4}(N)\left\lvert t-r\right\rvert^{2}$ (8.13)
@@
At this point we have upper bounds (8.15) for each term of (8.6). This means that, in order to verify (8.1), and to conclude the proof of Lemma 5.5, it is enough to show that for every $k\in\langle 5\rangle$
@@
[…]
@@
$\leqslant 2\mathcal{E}({H_{i}})+2\mathcal{E}({H_{j}})$
@@
$\mathcal{F}_{5}(N)\geqslant\mathcal{F}_{4}(N)\geqslant\mathcal{F}_{3}(N)\geqslant\mathcal{F}_{2}(N)=\mathcal{F}_{1}(N),$
@@
$=\frac{\mathcal{F}_{N}^{2}({H}_{i})\mathcal{F}_{N}^{2}({H}_{j})}{\mathcal{F}^{\rm opt}_{N}({H}_{i})\mathcal{F}^{\rm opt}_{N}({H}_{j})}\sum_{V=V_{\star}}^{V^{\star}}\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}_{V}(i,j)\\ \mathcal{V}(g_{\cup})=V\end{subarray}}\frac{1}{\mathcal{F}_{N}(g_{\cup})}$ (8.18)
@@
$\mathcal{E}(g_{1}\cap g_{2})\geqslant 1,\qquad\mathcal{E}(g_{3}\cap g_{4})\geqslant 1,.$
@@
$=\frac{\mathcal{F}_{N}(g_{1})\mathcal{F}_{N}(g_{2})}{\mathcal{F}_{N}(g_{1}\cap g_{2})}\frac{\mathcal{F}_{N}(g_{3})}{\mathcal{F}_{N}\bigl{(}\bigl{(}g_{1}\cup g_{2}\bigr{)}\cap g_{3}\bigr{)}}\frac{\mathcal{F}_{N}(g_{4})}{\mathcal{F}_{N}\bigl{(}\bigl{(}g_{1}\cup g_{2}\cup g_{3}\bigr{)}\cap g_{4}\bigr{)}}$ (8.19)
@@
Hence, from (8.19) we see that, for large enough $N$, denoting $g_{\star}=\bigl{(}g_{1}\cup g_{2}\cup g_{3}\bigr{)}\cap g_{4}\in{\tt CS}({H}_{j})$
$\mathcal{F}_{N}(g_{\cup})\geqslant\frac{\prod_{k=1}^{4}\mathcal{F}_{N}(g_{k})}{\mathcal{F}_{N}(g_{1}\cap g_{2})\mathcal{F}_{N}(g_{\star})}$ (8.20)
@@
$=\sum_{V=V_{\star}}^{V^{\star}}\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}_{V}(i,j)\\ \mathcal{V}(g_{\cup})=V\end{subarray}}\frac{\mathcal{F}_{N}(g_{1}\cap g_{2})}{\mathcal{F}^{\rm opt}_{N}({H}_{i})}\frac{\mathcal{F}_{N}(g_{\star})}{\mathcal{F}^{\rm opt}_{N}({H}_{j})}$ (8.22)
Using (3.10) for $g=g_{1}\cap g_{2}\in{\tt CS}(H_{i})$, we obtain for large enough $N$ that
@@
and using (3.10) for $g=g_{\star}\in{\tt CS}(H_{j})$ we obtain for large enough $N$ that
@@
$=4\sum_{V=V_{\star}}^{V^{\star}}\left(\frac{V!}{(V-\mathcal{V}(H_{i}))!\mathcal{A}(H_{i})}\right)^{2}\left(\frac{V!}{(V-\mathcal{V}(H_{j}))!\mathcal{A}(H_{j})}\right)^{2}=F^{\star}_{i,j},$
@@ Appendix A Auxiliary proofs
Proof of Proposition 2.6. We first show that in the setting considered the expected number of switching points of the alternating renewal process is finite. Denote for any given $N\in\mathbb{N}$ and starting state $x\in\{0,1\}$ the $i$-th (for $i\in\mathbb{N}$) switching point of the alternating renewal process $a_{N}(\cdot)$ by
@@
$\mathbb{E}\left\{\max\{n\colon\tau_{N,1,n}\leqslant T\}\right\}\leqslant\sum_{i=1}^{\infty}\int_{0}^{T}\frac{P^{i}t^{i-1}}{(i-1)!}\text{\rm d}t=e^{PT}-1$ (A.3)
@@
$\mathbb{E}\left\{\max\{n\colon\tau_{N,0,n}\leqslant T\}\right\}\leqslant\varrho_{N}(e^{PT}-1)$ (A.4)
@@
$=\sum_{{x\in\{0,1\}}}\mathbb{P}\left\{a_{N}(0)=x\right\}\sum_{{i=0}}^{\infty}\int_{0}^{r}\mathbb{P}\left\{\xi_{N,x,{i+1}}\in[r-t_{0},\,s-t_{0}]\right\}\text{\rm d}F_{\tau_{N,x,{i}}}(t_{0})$
Now applying the second and third assumption of Proposition 2.6, we conclude that
@@
+$\qquad+\mathbb{P}\left\{a_{N}(0)=0\right\}\left(P\varrho_{N}\left\lvert s-r\right\rvert+\sum_{i=1}^{\infty}\int_{0}^{r}P\left\lvert s-r\right\rvert\text{\rm d}F_{\tau_{N,0,i}}(t_{0})\right).$
Furthermore, using the first assumption of Proposition 2.6 for the probability $\mathbb{P}\left\{a_{N}(0)=1\right\}$, and combining (A.3) and (A.4), we obtain for large enough $N$ that
@@
$\mathbb{P}\left\{a_{N}(s)\not=a_{N}(r),\,a_{N}(t)\not=a_{N}(s)\right\}$
@@
Then applying the second assumption of Proposition 2.6 to the probability $\mathbb{P}\left\{\xi_{N,x,n+1}<\left\lvert t-r\right\rvert\right\}$, we have
@@
Proof of Lemma 3.4. So as to calculate $\mathcal{C}({H},{H^{\star}},g)$, we first choose $\mathcal{V}(g)$ vertices which belong to both graphs $g_{1}$ and $g_{2}$, and then divide the remaining $\mathcal{N}-\mathcal{V}(g)$ vertices between $g_{1}$ and $g_{2}$. To do this, we introduce some new notation. We define, for any ${\tt S}\subset\langle\mathcal{N}\rangle$, $V\leqslant\left\lvert{\tt S}\right\rvert$,
@@
$\mathcal{C}({H},{H^{\star}},g)$ $=\frac{1}{\mathcal{N}!}\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{\mathcal{N}}({H})\\ g_{2}\in{\tt G}_{\mathcal{N}}({H^{\star}})\end{subarray}}\mathbb{I}\{g_{1}\cap g_{2}\sim g\}$
$=\frac{1}{\mathcal{N}!}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{N}\rangle}(\mathcal{V}(g))}\sum_{\begin{subarray}{c}g^{\star}{\,\sqsubset\,}K_{\boldsymbol{v}}\\ g^{\star}\sim g\end{subarray}}\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{\mathcal{N}}({H})\\ g_{2}\in{\tt G}_{\mathcal{N}}({H^{\star}})\end{subarray}}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$
$=\frac{1}{\mathcal{N}!}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{N}\rangle}(\mathcal{V}(g))}\sum_{\boldsymbol{v}_{1}\in{\tt P}_{\langle N\rangle\setminus\boldsymbol{v}}(\mathcal{V}({H})-\mathcal{V}(g))}\sum_{\begin{subarray}{c}g^{\star}{\,\sqsubset\,}K_{\boldsymbol{v}}\\ g^{\star}\sim g\end{subarray}}\sum_{\begin{subarray}{c}g_{1}{\,\sqsubset\,}K_{\boldsymbol{v}\cup\boldsymbol{v}_{1}}\\ g_{1}\sim{H}\end{subarray}}\sum_{\begin{subarray}{c}g_{2}{\,\sqsubset\,}K_{\langle\mathcal{N}\rangle\setminus\boldsymbol{v}_{1}}\\ g_{2}\sim{H^{\star}}\end{subarray}}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$ (A.7)

### v2

@@ 8 Proof of Lemma 6
$\eta_{i,g,N}(s,t)$ $=\mathbb{I}\{g{\,{\subset}\,}G_{N}(t)\}-p^{\mathcal{E}({H}_{i})}(t)-\mathbb{I}\{g{\,{\subset}\,}G_{N}(s)\}+p^{\mathcal{E}({H}_{i})}(s).$ (8.3)
@@
$x_{k}=\begin{cases}s,\qquad k\in\{1,2\},\\ r,\qquad k\in\{3,4\},\end{cases}\qquad y_{k}=\begin{cases}t,\qquad k\in\{1,2\},\\ s,\qquad k\in\{3,4\}{,}\end{cases}\qquad i_{k}=\begin{cases}i,\qquad k\in\{1,2\},\\ j,\qquad k\in\{3,4\}.\end{cases}$
@@
$\Delta_{N,i,j}(r,s,t)=\frac{\mathcal{F}_{N}^{2}({H}_{i})\mathcal{F}_{N}^{2}({H}_{j})}{\mathcal{F}^{\rm opt}_{N}({H}_{i})\mathcal{F}^{\rm opt}_{N}({H}_{j})}\sum_{\boldsymbol{g}\in{{\tt G}_{N}(i,j)}}\mathbb{E}\left\{\prod_{k=1}^{4}\eta_{i_{k},g_{k},N}(x_{k},y_{k})\right\}{.}$ (8.6)
@@
$\Omega_{k}$ $=\bigl\{\mathbb{I}\{g_{k}{\,{\subset}\,}G_{N}(y_{k})\}\not=\mathbb{I}\{g_{k}{\,{\subset}\,}G_{N}(x_{k})\}\bigr\}.$ (8.7)
@@
Continuation of proof of Lemma 6. Combining the bounds presented in Lemma 9 with (8.8) and the evident bound $\eta_{i_{k},g_{k},N}(x_{k},y_{k})\leqslant 2$, we obtain that
$\hskip-34.14322pt\mathbb{E}\left\{\prod_{k=1}^{4}\eta_{i_{k},g_{k},N}(x_{k},y_{k})\right\}$
[…]
@@
[…]
@@
$16\,C_{0}\varrho_{N}^{\mathcal{E}(g_{1}\cup g_{2}\cup g_{3}\cup g_{4})}\left\lvert t-r\right\rvert^{2}{.}$
@@
$\prod_{k=1}^{4}\left(\varrho_{N}^{\mathcal{E}({H}_{i_{k}})}\mathcal{E}({H}_{i_{k}})p_{+}^{\mathcal{E}({H}_{i_{k}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert\right)\leqslant C_{1}F_{1}(N)\left\lvert t-r\right\rvert^{2}{,}$ (8.10)
@@
[…]
$\qquad\leqslant C_{2}F_{2}(N)\left\lvert t-r\right\rvert^{2}{,}$ (8.11)
@@
[…]
$\qquad\leqslant 4!\,C_{3}F_{3}(N)\left\lvert t-r\right\rvert^{2}{,}$ (8.12)
@@
$\frac{8}{3!}\sum_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}C_{0}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}})}\left\lvert t-r\right\rvert^{2}\varrho_{N}^{\mathcal{E}({H}_{i_{k_{4}}})}\mathcal{E}({H}_{i_{k_{4}}})p_{+}^{\mathcal{E}({H}_{i_{k_{4}}})-1}{\mathfrak{C}}\left\lvert t-r\right\rvert$
$\qquad\leqslant 4!\,C_{4}F_{4}(N)\left\lvert t-r\right\rvert^{2}{,}$ (8.13)
@@
At this point we have upper bounds (8.15) for each term of (8.6). This means that, in order to verify (8.1), and to conclude the proof of Lemma 6, it is enough to show that for every $k\in\langle 5\rangle$
@@
where $F^{*}_{i,j}$ and $C^{*}_{i,j}$ are defined in (5.5) and (5.6), respectively. The bound (8.17) is trivial, as $\mathfrak{C}\geqslant 1$, $p_{+}\geqslant 1$, $C_{0}\geqslant 1$ and $\mathcal{E}(H_{i_{k}})\geqslant 1$ for any $k\in\langle 4\rangle$ implying
$C_{1}$ $={\mathfrak{C}^{4}}\,T^{2}\,\mathcal{E}^{2}({H}_{i})\,\mathcal{E}^{2}({H}_{j})\,p_{+}^{2\mathcal{E}({H}_{i})+2\mathcal{E}({H}_{j})-4}\leqslant C^{\star}_{i,j},$
$C_{2}$ $=\frac{1}{3}\,C_{0}\,{\mathfrak{C}^{3}}\,T^{2}\max_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}\prod_{q=2}^{4}\left(\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}\right)$
$\leqslant\frac{1}{3}\,C_{0}\,{\mathfrak{C}^{3}}\,T^{2}\max_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}\prod_{q=1}^{4}\left(\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}\right)$
$=\frac{1}{3}\,C_{0}\,{\mathfrak{C}^{3}}\,T^{2}\mathcal{E}^{2}({H}_{i})\,\mathcal{E}^{2}({H}_{j})\,p_{+}^{2\mathcal{E}({H}_{i})+2\mathcal{E}({H}_{j})-4}\leqslant C_{i,j}^{\star},$
$C_{3}$ $=C_{0}\,{\mathfrak{C}^{2}}\,T\max_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}\prod_{q=3}^{4}\left(\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}\right)$
$\leqslant C_{0}\,{\mathfrak{C}^{2}}\,T\max_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}\prod_{q=1}^{4}\left(\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}\right)$
$=C_{0}\,{\mathfrak{C}^{2}}\,T\mathcal{E}^{2}({H}_{i})\,\mathcal{E}^{2}({H}_{j})\,p_{+}^{2\mathcal{E}({H}_{i})+2\mathcal{E}({H}_{j})-4}\leqslant C_{i,j}^{\star},$
$C_{4}$ $=\frac{4}{3}C_{0}\,{\mathfrak{C}}\,T\max_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}\mathcal{E}({H}_{i_{k_{4}}})\,p_{+}^{\mathcal{E}({H}_{i_{k_{4}}})-1}$
$\leqslant\frac{4}{3}C_{0}\,{\mathfrak{C}}\,T\max_{\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle}\prod_{q=1}^{4}\left(\mathcal{E}({H}_{i_{k_{q}}})p_{+}^{\mathcal{E}({H}_{i_{k_{q}}})-1}\right)$
$=\frac{4}{3}C_{0}\,{\mathfrak{C}}\,T\mathcal{E}^{2}({H}_{i})\,\mathcal{E}^{2}({H}_{j})\,p_{+}^{2\mathcal{E}({H}_{i})+2\mathcal{E}({H}_{j})-4}\leqslant C_{i,j}^{\star},$
$C_{5}$ $=16\,C_{0}\leqslant C_{i,j}^{\star}.$
This entails that we can consider only (8.16). For any set $\boldsymbol{g}\in{\tt G}_{N}(i,j)$ and any permutation $\{k_{1},k_{2},k_{3},k_{4}\}=\langle 4\rangle$ we have that
@@
$\leqslant 2\mathcal{E}({H_{i}})+2\mathcal{E}({H_{j}}){.}$
@@
${{F}_{5}(N)\geqslant{F}_{4}(N)\geqslant{F}_{3}(N)\geqslant{F}_{2}(N)={F}_{1}(N),}$
@@
$=\frac{\mathcal{F}_{N}^{2}({H}_{i})\mathcal{F}_{N}^{2}({H}_{j})}{\mathcal{F}^{\rm opt}_{N}({H}_{i})\mathcal{F}^{\rm opt}_{N}({H}_{j})}\sum_{V=V_{\star}}^{V^{\star}}\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}_{V}(i,j)\\ \mathcal{V}(g_{\cup})=V\end{subarray}}\frac{1}{\mathcal{F}_{N}(g_{\cup})}{.}$ (8.18)
@@
$\mathcal{E}(g_{1}\cap g_{2})\geqslant 1,\qquad\mathcal{E}(g_{3}\cap g_{4})\geqslant 1{.}$
@@
$=\frac{\mathcal{F}_{N}(g_{1})\mathcal{F}_{N}(g_{2})}{\mathcal{F}_{N}(g_{1}\cap g_{2})}\frac{\mathcal{F}_{N}(g_{3})}{\mathcal{F}_{N}\bigl(\bigl(g_{1}\cup g_{2}\bigr)\cap g_{3}\bigr)}\frac{\mathcal{F}_{N}(g_{4})}{\mathcal{F}_{N}\bigl(\bigl(g_{1}\cup g_{2}\cup g_{3}\bigr)\cap g_{4}\bigr)}{.}$ (8.19)
@@
Hence, from (8.19) we see that, for large enough $N$, denoting $g_{\star}=\bigl(g_{1}\cup g_{2}\cup g_{3}\bigr)\cap g_{4}$
$\mathcal{F}_{N}(g_{\cup})\geqslant\frac{\prod_{k=1}^{4}\mathcal{F}_{N}(g_{k})}{\mathcal{F}_{N}(g_{1}\cap g_{2})\mathcal{F}_{N}(g_{\star})}{.}$ (8.20)
@@
$=\sum_{V=V_{\star}}^{V^{\star}}\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}_{V}(i,j)\\ \mathcal{V}(g_{\cup})=V\end{subarray}}\frac{\mathcal{F}_{N}(g_{1}\cap g_{2})}{\mathcal{F}^{\rm opt}_{N}({H}_{i})}\frac{\mathcal{F}_{N}(g_{\star})}{\mathcal{F}^{\rm opt}_{N}({H}_{j})}{.}$ (8.22)
Using (3.9) for $g=\pi(g_{1}\cap g_{2})\in{\tt CS}(H_{i})$ we obtain for large enough $N$ that
@@
and using (3.9) for $g=\pi(g_{\star})\in{\tt CS}(H_{j})$ we obtain for large enough $N$ that
@@
$=4\sum_{V=V_{\star}}^{V^{\star}}\left(\frac{V!}{(V-\mathcal{V}(H_{i}))!\mathcal{A}(H_{i})}\right)^{2}\left(\frac{V!}{(V-\mathcal{V}(H_{j}))!\mathcal{A}(H_{j})}\right)^{2}$
$=F^{\star}_{i,j},$
@@ Appendix A Auxiliary proofs
Proof of Proposition 1. We first show that in the setting considered the expected number of switching points of the alternating renewal process is finite. Denote for any given $N\in\mathbb{N}$ and starting state $x\in\{0,1\}$ the $i$-th (for $i\in\mathbb{N}$) switching point of the alternating renewal process $a_{N}(\cdot)$ by
@@
$\mathbb{E}\left\{\max\{n\colon\tau_{N,1,n}\leqslant T\}\right\}\leqslant\sum_{i=1}^{\infty}\int_{0}^{T}\frac{P^{i}t^{i-1}}{(i-1)!}\text{\rm d}t=e^{PT}-1{.}$ (A.3)
@@
$\mathbb{E}\left\{\max\{n\colon\tau_{N,0,n}\leqslant T\}\right\}\leqslant\varrho_{N}(e^{PT}-1){.}$ (A.4)
@@
$=\sum_{{x\in\{0,1\}}}\mathbb{P}\left\{a_{N}(0)=x\right\}\sum_{{i=0}}^{\infty}\int_{0}^{r}\mathbb{P}\left\{\xi_{N,x,{i+1}}\in[r-t_{0},\,s-t_{0}]\right\}\text{\rm d}F_{\tau_{N,x,{i}}}(t_{0}){.}$
Now applying the second and third assumption of Proposition 1, we conclude that
@@
$\qquad+\mathbb{P}\left\{a_{N}(0)=0\right\}\left(P\varrho_{N}\left\lvert s-r\right\rvert+\sum_{i=1}^{\infty}\int_{0}^{r}P\left\lvert s-r\right\rvert\text{\rm d}F_{\tau_{N,0,i}}(t_{0})\right).$
Furthermore, using the first assumption of Proposition 1 for the probability $\mathbb{P}\left\{a_{N}(0)=1\right\}$, and combining (A.3) and (A.4), we obtain for large enough $N$ that
@@
$\hskip-17.07164pt\mathbb{P}\left\{a_{N}(s)\not=a_{N}(r),\,a_{N}(t)\not=a_{N}(s)\right\}$
@@
Then applying the second assumption of Proposition 1 to the probability $\mathbb{P}\left\{\xi_{N,x,n+1}<\left\lvert t-r\right\rvert\right\}$, we have
@@
Proof of Lemma 2. In order to keep our notation compact, in this proof we locally abbreviate $\mathcal{N}(H,H^{\star},g)$ to $\mathcal{N}$. So as to calculate $\mathcal{C}({H},{H^{\star}},g)$, we first choose $\mathcal{V}(g)$ vertices which belong to both graphs $g_{1}$ and $g_{2}$, and then divide the remaining $\mathcal{N}-\mathcal{V}(g)$ vertices between $g_{1}$ and $g_{2}$. To do this, we introduce some new notation. We define, for any ${\tt S}\subset\langle\mathcal{N}\rangle$, $V\leqslant\left\lvert{\tt S}\right\rvert$,
@@
$\mathcal{C}({H},{H^{\star}},g)$ $=\frac{1}{\mathcal{N}!}\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{\mathcal{N}}({H})\\ g_{2}\in{\tt G}_{\mathcal{N}}({H^{\star}})\end{subarray}}\mathbb{I}\{{\pi(g_{1}\cap g_{2})=g}\}$
$=\frac{1}{\mathcal{N}!}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{N}\rangle}(\mathcal{V}(g))}\sum_{\begin{subarray}{c}g^{\star}{\,{\subset}\,}K_{\boldsymbol{v}}\\ {\pi(g^{\star})=g}\end{subarray}}\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{\mathcal{N}}({H})\\ g_{2}\in{\tt G}_{\mathcal{N}}({H^{\star}})\end{subarray}}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$
$=\frac{1}{\mathcal{N}!}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{N}\rangle}(\mathcal{V}(g))}\sum_{\boldsymbol{v}_{1}\in{\tt P}_{\langle N\rangle\setminus\boldsymbol{v}}(\mathcal{V}({H})-\mathcal{V}(g))}\sum_{\begin{subarray}{c}g^{\star}{\,{\subset}\,}K_{\boldsymbol{v}}\\ {\pi(g^{\star})=g}\end{subarray}}\sum_{\begin{subarray}{c}g_{1}{\,{\subset}\,}K_{\boldsymbol{v}\cup\boldsymbol{v}_{1}}\\ g_{1}\sim{H}\end{subarray}}\sum_{\begin{subarray}{c}g_{2}{\,{\subset}\,}K_{\langle\mathcal{N}\rangle\setminus\boldsymbol{v}_{1}}\\ g_{2}\sim{H^{\star}}\end{subarray}}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}{.}$ (A.7)